Q4. Find the domain and graph the functions. Also tell whether they are

one-one or onto.

a. √|π‘₯|

b. 1βˆ’2π‘₯βˆ’π‘₯2


1
Expert's answer
2021-04-15T07:40:05-0400

a. "y = \sqrt {|x|}"

Solution:

square root expression must be non-negative, so "|x| \ge 0" - performed for any "x \in \left( { - \infty ; + \infty } \right)". Then "D(y):\,x \in \left( { - \infty ; + \infty } \right)" .

Plot the function:



Function is on to if "\forall y \in Y\exists x \in X:\,f(x) = y" , so, we have onto function.

b. "y = 1 - 2x - {x^2}"

Solution:

There is no restriction for the variable x, so "D(y):\,x \in \left( { - \infty ; + \infty } \right)" .

The function graph is a parabola. Let's find the coordinates of the vertex:

"{x_0} = - \frac{b}{{2a}} = - \frac{{ - 2}}{{ - 2}} = - 1 \Rightarrow {y_0} = 1 + 2 - 1 = 2"

Find the zeros of the function:

"{x^2} + 2x - 1 = 0"

"D = 4 + 4 = 8"

"{x_1} = \frac{{ - 2 - \sqrt 8 }}{2} = - 1 - \sqrt 2 ,\,\,{x_2} = - 1 + \sqrt 2"

We have the graph:




similarly, we have onto function.